# Random Lévy Looptrees and Lévy Maps

### v2

@@ 3.2.   A volume estimate in the looptree
Notice that we give them all the same drift $\beta$ for the continuous part. Informally, in the infinite looptree coded by $(\overleftarrow{X},\overrightarrow{X})$ with a distinguished path of loops by coded $A_{0}\cup B_{0}$, when reading this path starting from 00, the jumps of process $X^{\prime}$ code the total length of the loops on this path, those of $X^{R}$ code their right length, those of $\widetilde{X}$ the shortest length among left and right, so related to the looptree distance, and finally those of $\sigma$ the product of the left and right lengths, which is related to a Brownian bridge on the loop.
@@
$\psi^{\prime}(\lambda)-\beta\lambda\quad\text{for }X^{\prime},\qquad\widetilde{\psi}(\lambda)\quad\text{for }X^{R},\qquad\widetilde{\psi}(\lambda/2)+\beta\lambda/2\quad\text{for }\widetilde{X},\quad\text{and}\quad\phi(\lambda)+\beta\lambda\quad\text{for }\sigma,$
@@
$-\frac{1}{t}\log\mathop{}\!\mathbb{E}\Bigl{[}\exp\Bigl{(}-\sum_{t_{i}\leq t}x_{i}\Bigr{)}\Bigr{]}=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})x\pi(\mathop{}\!\mathrm{d}x)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\lambda}\int_{0}^{\infty}(\operatorname{e}^{-\lambda x}-1+\lambda x)\pi(\mathop{}\!\mathrm{d}x)=\psi^{\prime}(\lambda)-2\beta\lambda.$
For $X^{R}$, we have:
$-\frac{1}{t}\log\mathop{}\!\mathbb{E}\Bigl{[}\exp\Bigl{(}-\sum_{t_{i}\leq t}u_{i}x_{i}\Bigr{)}\Bigr{]}=\int_{0}^{\infty}x\pi(\mathop{}\!\mathrm{d}x)\int_{0}^{1}(1-\operatorname{e}^{-\lambda ux})\mathop{}\!\mathrm{d}u=\frac{1}{\lambda}\int_{0}^{\infty}\pi(\mathop{}\!\mathrm{d}x)(\lambda x-1+\operatorname{e}^{-\lambda x})=\widetilde{\psi}(\lambda)-\beta\lambda.$
@@
$-\frac{1}{t}\log\mathop{}\!\mathbb{E}\Bigl{[}\exp\Bigl{(}-\sum_{t_{i}\leq t}x_{i}\min\{u_{i},1-u_{i}\}\Bigr{)}\Bigr{]}=\int_{0}^{\infty}x\pi(\mathop{}\!\mathrm{d}x)\int_{0}^{1/2}2\mathop{}\!\mathrm{d}v(1-\operatorname{e}^{-\lambda vx})=\widetilde{\psi}\Bigl{(}\frac{\lambda}{2}\Bigr{)}-\frac{\beta\lambda}{2}.$
@@
$\mathop{}\!\mathbb{P}\Bigl{(}\sum_{t_{i}\leq\tau_{\varepsilon}-}x_{i}\mathbf{1}_{\{v_{i}\geq\frac{1}{2K_{\varepsilon}}\}}\geq\varepsilon^{1-\delta}\bigm{|}(t_{i},y_{i})_{i\in I}\Bigr{)}$ $\leq\mathop{}\!\mathbb{P}\Bigl{(}\sum_{t_{i}\leq\tau_{\varepsilon}-}\frac{1}{v_{i}}y_{i}\mathbf{1}_{\{v_{i}\geq\frac{1}{2K_{\varepsilon}}\}}\geq\varepsilon^{-\delta}\sum_{t_{i}\leq\tau_{\varepsilon}-}y_{i}\bigm{|}(t_{i},y_{i})_{i\in I}\Bigr{)}$
$\leq 2\varepsilon^{\delta}\log K_{\varepsilon}$
$\leq 2\varepsilon^{\delta-\delta^{\prime}}.$
@@
$\mathop{}\!\mathbb{P}\Bigl{(}\sum_{t_{i}\leq t}x_{i}\mathbf{1}_{\{v_{i}<\frac{1}{2K_{\varepsilon}}\}}>x\Bigr{)}=\mathop{}\!\mathbb{P}(1-\operatorname{e}^{X^{\prime}_{t/K_{\varepsilon}}/x}>1-\operatorname{e}^{-1})\leq\frac{\operatorname{e}}{\operatorname{e}-1}\bigl{(}1-\mathop{}\!\mathbb{E}\bigl{[}\operatorname{e}^{-X^{\prime}_{t/K_{\varepsilon}}/x}\bigr{]}\bigr{)}\leq\frac{\operatorname{e}}{\operatorname{e}-1}\frac{t}{K_{\varepsilon}}\psi^{\prime}(1/x).$
On the other hand, since $\widetilde{X}$ has Laplace exponent $\widetilde{\psi}(\cdot/2)$,
@@
which is a convergent series as we wanted. This implies that almost surely, for every $n$ large enough, we have $0^{+}_{2^{-n}}\geq\psi(2^{n(1-\delta)})^{-1}$ in the infinite volume model. By local absolute continuity, we deduce that almost surely, for every $n$ large enough, we have $U^{+}_{2^{-n}}-U\leq\psi(2^{n(1-\delta)})^{-1}$ in the excursion model.
@@
Recall from the previous proof of that $B(u,r)$ denotes the closed ball centred at $U$ and with radius $r$ in $\mathscr{L}(X^{\mathrm{ex}})$, seen as a subset of $[0,1]$ and that ||$\cdot$|| ⋅ | denotes the Lebesgue measure.
@@
Fix $u\in[0,1]$. Combining Remark 2.2 and Remark 2.3 we infer the following: there exist discrete Łukasiewicz paths, obtained by cyclicly shifting a bridge with exchangeable increments, which converge in distribution to $X^{\mathrm{ex}}$ once properly rescaled, such that if $d^{n}$ denotes the corresponding looptree distance, normalised both in time and space, then $(d^{n}(u,t))_{t\in[0,1]}$ converges to $(d_{\mathscr{L}(X^{\mathrm{ex}})}(u,t))_{t\in[0,1]}$ for the uniform topology. In particular, the processes $(|B(u,r)\cap[0,u]|)_{r\geq 0}$ and $(|B(1-u,r)\cap[1-u,1]|)_{r\geq 0}$ are the limits in law of the analogous processes for these rescaled discrete looptrees. We claim that these discrete volume processes for any given $n$ have the same distribution as in fact the discrete looptrees are almost invariant under mirror symmetry.
Refer to captionRefer to captionRefer to caption
@@ 3.3.   On distances in the map
Suppose next that we are in the first case in (3.6), but now that there exists $r\in(s,t)$ with $X^{\mathrm{ex}}_{r}=X^{\mathrm{ex}}_{s}=X^{\mathrm{ex}}_{t}$. Then by uniqueness of the local minima, neither $s$ nor $t$ are local minima. In addition, as recalled in the proof of Lemma 3.3, we have $\Delta X^{\mathrm{ex}}_{s}=0$ in this case since otherwise the reversed process at time $r$ would may a weak record by a jump. This implies that for every $\varepsilon>0$, there exist $s^{\prime}\in(s-\varepsilon,s)$ and $t^{\prime}\in(t,t+\varepsilon)$ which satisfy $X^{\mathrm{ex}}_{s^{\prime}}=X^{\mathrm{ex}}_{t^{\prime}}<X^{\mathrm{ex}}_{r}$ for all $r\in(s^{\prime},t^{\prime})$ and we conclude from the preceding case.
@@
Let us next continue with a geometric lemma inspired by [LGM11, Lemma 14] which will allow us to include balls in the map in some well-chosen intervals analogously to Lemma 3.4 for looptrees. Recall the sets $A_{u}$ and $B_{u}$ defined in (3.1) and (3.2) respectively.
@@
We shall assume throughout the proof of that both $Z_{a}<Z_{u}$ and $Z_{b}<Z_{u}$ otherwise the claim is trivial.
@@
We infer from the convergence $X^{n}\to X^{\mathrm{ex}}$ in the Skorokhod $J_{1}$ topology that there exist times $s_{n}<a_{n}<u_{n}<b_{n}$ in $\mathbb{N}/n$ which converge to $s$, $a$, $u$, $b$ respectively and such that $X^{n}_{a_{n}}=\min_{[s_{n},a_{n}]}X^{n}>X^{n}_{b_{n}}=\min_{[s_{n},b_{n}]}X^{n}>X^{n}_{s_{n}-1/n}$. This implies that $a_{n}$ and $b_{n}$ both belong to the same cycle between two consecutive ancestors of $u_{n}$, with $a_{n}$ on the left and $b_{n}$ on the right and we infer from the first part of the proof of that the graph distance in the rescaled map between $u_{n}$ and any vertex $t_{n}$ which lies outside $[a_{n},b_{n}]$ is larger than or equal to $\min\{Z^{n}_{u_{n}}-Z^{n}_{a_{n}},Z^{n}_{u_{n}}-Z^{n}_{b_{n}}\}$. We finally can pass to the limit along the corresponding subsequence, using that $Z^{n}\to Z$ for the uniform topology.
@@ 3.4.   A technical lemma with Brownian motions
In the case of stable Lévy processes, Le Gall & Miermont [LGM11, Lemma 15] were able to find such a pair $(a_{r},b_{r})\in A_{0}$, namely that codes an ancestor of the distinguished point 00. Their argument can be extended when the upper and lower exponents $\boldsymbol{\gamma}$ and $\boldsymbol{\eta}$ coincide, or when the exponent $\boldsymbol{\delta}\in[1,\boldsymbol{\gamma}]$ introduced Duquesne [Duq12] as $\boldsymbol{\delta}\coloneqq\sup\{c>0\colon\exists K\in(0,\infty)\text{ such that }K\,\mu^{-c}\psi(\mu)\leq\lambda^{-c}\psi(\lambda),\text{ for all }1\leq\mu\leq\lambda\}$ is not equal to $1$. However we were not able to extend it when $\boldsymbol{\delta}=1$. We therefore argue in a general setting, conditionally on the underlying looptree, and relying only on the properties of the Brownian labels.
@@
  • or there exist $i\geq 1$ with $g_{i}<r^{2(1-\varepsilon)}$ and $s\in[0,\min\{\delta^{L}_{i},r^{2(1-\varepsilon)}\}]$ and $t\in[0,\min\{\delta^{R}_{i},r^{2(1-\varepsilon)}\}]$ such that both $W^{L,i}_{s}\leq-r$ and $W^{R,i}_{t}\leq-r$.
The key point in this statement is that in the second case, both bridges reach $-r$ in the same interval of $\mathcal{C}^{c}\cap[0,r^{2(1-\varepsilon)}]$. The idea of the proof of is quite intuitive: first we know by the law of iterated logarithm that $W$ reaches values much less than $-r$ on $[0,r^{2(1-\varepsilon)}]$ so, if the first alternative fails, then there exists an interval $(g_{i},d_{i})$ on which $W$ goes much below $-r$, while starting and ending above this value. Then this interval cannot be too small for the Brownian motion to fluctuate enough. Therefore, if one resamples two Brownian bridges with the same duration $\delta_{i}$ and the same endpoints $W_{g_{i}}$ and $W_{d_{i}}$, then they have a reasonable chance to both go below $-r$ in this interval. Actually the Brownian bridges we sample have a longer duration, so intuitively, they fluctuate even more and have an even greater chance to reach $-r$. However turning this discussion into a rigorous proof of is quite challenging, and the proof of Lemma 3.11 is deferred to Section 3.5.
@@
Recall the Poisson random measure $\mathscr{N}$ in (3.5) and the four subordinators defined from it. By construction of the looptree, the successive ancestors of 00 in the infinite looptree, started from 00, are coded by the negative of the first element of pairs in $A_{0}$ (defined in (3.1)) and correspond to the times of weak records of $\overleftarrow{X}$. Further, by construction of the label process, the labels of these ancestors form a subordinate Brownian motion, precisely $W_{\sigma(\cdot)}$, where $W$ is a standard Brownian motion and $\sigma$ is the last subordinator in Lemma 3.5. We shall apply Lemma 3.11 to the closed set $\mathcal{C}$ given by the range of $\sigma$.
@@ 3.5.   Proof of the technical lemma
$\mathop{}\!\mathbb{P}(\tau_{-r\log^{10K_{r}}(1/r)}>r^{2(1-1/\log\log(1/r))})$ $=\mathop{}\!\mathbb{P}(\inf\{W_{t},t\in[0,r^{2(1-1/\log\log(1/r))}]\}\geq-r\log^{10K_{r}}(1/r))$
$=\mathop{}\!\mathbb{P}(|W_{r^{2(1-1/\log\log(1/r))}}|\leq r\log^{10K_{r}}(1/r))$
$=\mathop{}\!\mathbb{P}(|W_{1}|\leq r^{1/\log\log(1/r)}\log^{10K_{r}}(1/r)).$
@@
For the last claim, by symmetry, for any given $k$, the probability that $W$ reaches $-r\log^{10k}(1/r)$ before $r\log^{10k}(1/r)$ equals $1/2$ so if these events were independent, then this would follow e.g. from Hoeffding’s inequality. They are not independent, however by the previous bound, we may assume that at the first time $W$ reaches $-r\log^{10k}(1/r)$, it has not reached $r\log^{10(k+1)}(1/r)$ yet, and by the strong Markov property, the probability starting from this position to reach $-r\log^{10(k+1)}(1/r)$ before $r\log^{10(k+1)}(1/r)$ is (slightly) larger than $1/2$. Formally, one can e.g. apply the Azuma inequality to the martingale whose increments are
$\Delta M_{k}=\mathbf{1}_{\{\tau_{-r\log^{10(k+1)}(1/r)}<\tau_{r\log^{10(k+1)}(1/r)}\}}-\mathop{}\!\mathbb{P}(\tau_{-r\log^{10(k+1)}(1/r)}<\tau_{r\log^{10(k+1)}(1/r)}\mid(W_{t},t\leq\tau_{-r\log^{10k}(1/r)})),$
@@
(#$\mathop{}\!\mathbb{P}\Bigl{(}\#\{k\leq K_{r}\colon\tau_{-r\log^{10(k+1)}(1/r)}<\tau_{r\log^{10(k+1)}(1/r)}\}<\frac{K_{r}}{4}\Bigr{)}$ $=\mathop{}\!\mathbb{P}\Bigl{(}\sum_{k=1}^{K_{r}}\Bigl{(}\mathbf{1}_{\{\tau_{-r\log^{10(k+1)}(1/r)}<\tau_{r\log^{10(k+1)}(1/r)}\}}-\frac{1}{2}\Bigr{)}<-\frac{K_{r}}{4}\Bigr{)}$
$\leq\mathop{}\!\mathbb{P}\Bigl{(}\sum_{k=1}^{K_{r}}\Delta M_{k}<-\frac{K_{r}}{4}\Bigr{)}$
$\leq\exp(-cK_{r}),$
@@
Similarly, for any $z>0$, the law of the time-reversed path $(z+W_{(\tau_{-z}-t)^{+}})_{t\geq 0}$ is that of a three-dimensional Bessel process started from 00, or equivalently the Euclidean norm of a three-dimensional Brownian motion, stopped at the last passage at $z$. The probability that a Bessel process does not reach $z/2$ before time $z^{2}$ is bounded below by some constant $c>0$ independent of $z$. We conclude as above, that the proportion of the number of indices $k\leq K_{r}$ such that this occurs for $z=r\log^{10k}(1/r)$ is bounded away from 00. ∎
@@
by a quantity that only depends on $r$ and whose sum over all values of $r\in\{2^{-n},n\geq 1\}$ is finite, such as $\log^{-2}(1/r)$. We then conclude as in the previous proof of from the Borel–Cantelli lemma.
Step 1: Proof of for Brownian motions. Let us first prove the claim when the Brownian bridges $W^{L}$ and $W^{R}$ are replaced by Brownian motions. Precisely, for every $k\leq K_{r}$, conditionally on $\Delta_{k}$, $\Delta_{k}^{L}$, $\Delta_{k}^{R}$, and $W_{G_{k}}$, let $X^{L,k}=(X^{L,k}_{t};t\in[0,\Delta_{k}^{L}])$ and $X^{R,k}=(X^{R,k}_{t};t\in[0,\Delta_{k}^{L}])$ be two independent Brownian motions started from $X^{L,k}_{0}=X^{R,k}_{0}=W_{G_{k}}>-r$.

### v3

@@ 4.2.   A volume estimate in the looptree
Notice that we give them all the same drift $\beta$ for the continuous part. Informally, in the infinite looptree coded by $(\overleftarrow{X},\overrightarrow{X})$ with a distinguished path of loops coded by $A_{0}\cup B_{0}$, when reading this path starting from 00, the jumps of process $X^{\prime}$ code the total length of the loops on this path, those of $X^{R}$ code their right length, those of $\widetilde{X}$ the shortest length among left and right, so related to the looptree distance, and finally those of $\sigma$ the product of the left and right lengths, which is related to a Brownian bridge on the loop.
@@
$\psi^{\prime}(\lambda)-\beta\lambda\,\,\,\text{for }X^{\prime},\,\,\,\widetilde{\psi}(\lambda)\,\,\,\text{for }X^{R},\,\,\,\widetilde{\psi}(\lambda/2)+\beta\lambda/2\,\,\,\text{for }\widetilde{X},\,\,\,\text{and}\,\,\,\phi(\lambda)+\beta\lambda\,\,\text{for }\sigma,$
@@
$-\frac{1}{t}\log\mathop{}\!\mathbb{E}\Bigl{[}\operatorname{e}^{-\sum_{t_{i}\leq t}x_{i}}\Bigr{]}=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})x\pi(\mathop{}\!\mathrm{d}x)=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}\lambda}\int_{0}^{\infty}(\operatorname{e}^{-\lambda x}-1+\lambda x)\pi(\mathop{}\!\mathrm{d}x),$
which is equal to $\psi^{\prime}(\lambda)-2\beta\lambda$. For $X^{R}$, we have:
$-\frac{1}{t}\log\mathop{}\!\mathbb{E}\Bigl{[}\operatorname{e}^{-\sum_{t_{i}\leq t}u_{i}x_{i}}\Bigr{]}=\int_{0}^{\infty}x\pi(\mathop{}\!\mathrm{d}x)\int_{0}^{1}(1-\operatorname{e}^{-\lambda ux})\mathop{}\!\mathrm{d}u=\frac{1}{\lambda}\int_{0}^{\infty}\pi(\mathop{}\!\mathrm{d}x)(\lambda x-1+\operatorname{e}^{-\lambda x}),$
which is equal to $\widetilde{\psi}(\lambda)-\beta\lambda$.
@@
$-\frac{1}{t}\log\mathop{}\!\mathbb{E}\Bigl{[}\exp\Bigl{(}-\sum_{t_{i}\leq t}x_{i}\min\{u_{i},1-u_{i}\}\Bigr{)}\Bigr{]}=\int_{0}^{\infty}x\pi(\mathop{}\!\mathrm{d}x)\int_{0}^{1/2}2\mathop{}\!\mathrm{d}v(1-\operatorname{e}^{-\lambda vx}),$
which is equal to $\widetilde{\psi}(\frac{\lambda}{2})-\frac{\beta\lambda}{2}$.
@@
$\mathop{}\!\mathbb{P}\Bigl{(}\sum_{t_{i}\leq\tau_{\varepsilon}-}x_{i}\mathbf{1}_{\{v_{i}\geq\frac{1}{2K_{\varepsilon}}\}}\geq\varepsilon^{1-\delta}\bigm{|}(t_{i},y_{i})_{i\in I}\Bigr{)}$
$\qquad\qquad\leq\mathop{}\!\mathbb{P}\Bigl{(}\sum_{t_{i}\leq\tau_{\varepsilon}-}\frac{1}{v_{i}}y_{i}\mathbf{1}_{\{v_{i}\geq\frac{1}{2K_{\varepsilon}}\}}\geq\varepsilon^{-\delta}\sum_{t_{i}\leq\tau_{\varepsilon}-}y_{i}\bigm{|}(t_{i},y_{i})_{i\in I}\Bigr{)}\leq 2\varepsilon^{\delta}\log K_{\varepsilon}\leq 2\varepsilon^{\delta-\delta^{\prime}}.$
@@
$\mathop{}\!\mathbb{P}\Bigl{(}\sum_{t_{i}\leq t}x_{i}\mathbf{1}_{\{v_{i}<\frac{1}{2K_{\varepsilon}}\}}>x\Bigr{)}=\mathop{}\!\mathbb{P}(1-\operatorname{e}^{X^{\prime}_{t/K_{\varepsilon}}/x}>1-\operatorname{e}^{-1})\leq\frac{\operatorname{e}}{\operatorname{e}-1}\bigl{(}1-\mathop{}\!\mathbb{E}\bigl{[}\operatorname{e}^{-X^{\prime}_{t/K_{\varepsilon}}/x}\bigr{]}\bigr{)},$
which is at most $\frac{\operatorname{e}}{\operatorname{e}-1}\frac{t}{K_{\varepsilon}}\psi^{\prime}(1/x)$. On the other hand, since $\widetilde{X}$ has Laplace exponent $\widetilde{\psi}(\cdot/2)$,
@@
which is a convergent series as we wanted. This implies that almost surely, for every $n$ large enough, we have $0^{+}_{2^{-n}}\leq\psi(2^{n(1-\delta)})^{-1}$ in the infinite volume model. By local absolute continuity, we deduce that almost surely, for every $n$ large enough, we have $U^{+}_{2^{-n}}-U\leq\psi(2^{n(1-\delta)})^{-1}$ in the excursion model.
@@
Recall from the previous proof that $B(u,r)$ denotes the closed ball centred at $U$ and with radius $r$ in $\mathscr{L}(X^{\mathrm{ex}})$, seen as a subset of $[0,1]$ and that ||$\cdot$|| ⋅ | denotes the Lebesgue measure.
@@
Fix $u\in[0,1]$. We infer from Remark 2.2 that there exist discrete Łukasiewicz paths, obtained by cyclicly shifting a bridge with exchangeable increments, which converge in distribution to $X^{\mathrm{ex}}$ once properly rescaled, such that if $d^{n}$ denotes the corresponding looptree distance, normalised both in time and space, then $(d^{n}(u,t))_{t\in[0,1]}$ converges to $(d_{\mathscr{L}(X^{\mathrm{ex}})}(u,t))_{t\in[0,1]}$ for the uniform topology. In particular, the processes $(|B(u,r)\cap[0,u]|)_{r\geq 0}$ and $(|B(1-u,r)\cap[1-u,1]|)_{r\geq 0}$ are the limits in law of the analogous processes for these rescaled discrete looptrees. We claim that these discrete volume processes for any given $n$ have the same distribution as in fact the discrete looptrees are almost invariant under mirror symmetry.
@@ 4.3.   On distances in the map
Suppose next that we are in the first case in (4.6), but now that there exists $r\in(s,t)$ with $X^{\mathrm{ex}}_{r}=X^{\mathrm{ex}}_{s}=X^{\mathrm{ex}}_{t}$. Then by uniqueness of the local minima, neither $s$ nor $t$ are local minima. Since in addition $\Delta X^{\mathrm{ex}}_{s}=0$, this then implies that for every $\varepsilon>0$, there exist $s^{\prime}\in(s-\varepsilon,s)$ and $t^{\prime}\in(t,t+\varepsilon)$ which satisfy $X^{\mathrm{ex}}_{s^{\prime}}=X^{\mathrm{ex}}_{t^{\prime}}<X^{\mathrm{ex}}_{r}$ for all $r\in(s^{\prime},t^{\prime})$ and we conclude from the preceding case.
@@
Let us next continue with a geometric lemma inspired by [LGM11, Lemma 14] which will allow us to include balls in the map in some well-chosen intervals analogously to Lemma 4.4 for looptrees. Recall the sets $A_{u}$ and $B_{u}$ defined by (4.1) and (4.2) respectively.
@@
We shall assume throughout the proof that both $Z_{a}<Z_{u}$ and $Z_{b}<Z_{u}$ otherwise the claim is trivial.
@@
We infer from the convergence $X^{n}\to X^{\mathrm{ex}}$ in the Skorokhod $J_{1}$ topology that there exist times $s_{n}<a_{n}<u_{n}<b_{n}$ in $\mathbb{N}/n$ which converge to $s$, $a$, $u$, $b$ respectively and such that $X^{n}_{a_{n}}=\min_{[s_{n},a_{n}]}X^{n}>X^{n}_{b_{n}}=\min_{[s_{n},b_{n}]}X^{n}>X^{n}_{s_{n}-1/n}$. This implies that $a_{n}$ and $b_{n}$ both belong to the same cycle between two consecutive ancestors of $u_{n}$, with $a_{n}$ on the left and $b_{n}$ on the right and we infer from the first part of the proof that the graph distance in the rescaled map between $u_{n}$ and any vertex $t_{n}$ which lies outside $[a_{n},b_{n}]$ is larger than or equal to $\min\{Z^{n}_{u_{n}}-Z^{n}_{a_{n}},Z^{n}_{u_{n}}-Z^{n}_{b_{n}}\}$. We finally can pass to the limit along the corresponding subsequence, using that $Z^{n}\to Z$ for the uniform topology.
@@ 4.4.   A technical lemma with Brownian motions
In the case of stable Lévy processes, Le Gall & Miermont [LGM11, Lemma 15] were able to find such a pair $(a_{r},b_{r})\in A_{0}$, namely one that codes an ancestor of the distinguished point 00. Their argument can be extended when the upper and lower exponents $\boldsymbol{\gamma}$ and $\boldsymbol{\eta}$ coincide, or when the exponent $\boldsymbol{\delta}\in[1,\boldsymbol{\gamma}]$ introduced by Duquesne [Duq12] as $\boldsymbol{\delta}\coloneqq\sup\{c>0\colon\exists K\in(0,\infty)\text{ such that }K\,\mu^{-c}\psi(\mu)\leq\lambda^{-c}\psi(\lambda),\text{ for all }1\leq\mu\leq\lambda\}$ is not equal to $1$. However we were not able to extend it when $\boldsymbol{\delta}=1$. We therefore argue in a general setting, conditionally on the underlying looptree, and relying only on the properties of the Brownian labels.
@@
  • or there exist $i\geq 1$ such that $g_{i}<r^{2(1-\varepsilon)}$, and $s\in[0,\min\{\delta^{L}_{i},r^{2(1-\varepsilon)}\}]$, and also finally $t\in[0,\min\{\delta^{R}_{i},r^{2(1-\varepsilon)}\}]$ such that both $W^{L,i}_{s}\leq-r$ and $W^{R,i}_{t}\leq-r$.
The key point in this statement is that in the second case, both bridges reach $-r$ in the same interval of $\mathcal{C}^{c}\cap[0,r^{2(1-\varepsilon)}]$. The idea of the proof is quite intuitive: first we know by the law of iterated logarithm that $W$ reaches values much less than $-r$ on $[0,r^{2(1-\varepsilon)}]$ so, if the first alternative fails, then there exists an interval $(g_{i},d_{i})$ on which $W$ goes much below $-r$, while starting and ending above this value. Then this interval cannot be too small for the Brownian motion to fluctuate enough. Therefore, if one resamples two Brownian bridges with the same duration $\delta_{i}$ and the same endpoints $W_{g_{i}}$ and $W_{d_{i}}$, then they have a reasonable chance to both go below $-r$ in this interval. Actually the Brownian bridges we sample have a longer duration, so intuitively, they fluctuate even more and have an even greater chance to reach $-r$. However turning this discussion into a rigorous proof is quite challenging, and the proof of Lemma 4.11 is deferred to Section 4.5.
@@
Recall the Poisson random measure $\mathscr{N}$ in (4.5) and the four subordinators defined from it. By construction of the looptree, the successive ancestors of 00 in the infinite looptree, started from 00, are coded by the negative of the first element of pairs in $A_{0}$ (defined by (4.1)) and correspond to the times of weak records of $\overleftarrow{X}$. Further, by construction of the label process, the labels of these ancestors form a subordinate Brownian motion, precisely $W_{\sigma(\cdot)}$, where $W$ is a standard Brownian motion and $\sigma$ is the last subordinator in Lemma 4.5. We shall apply Lemma 4.11 to the closed set $\mathcal{C}$ given by the range of $\sigma$.
@@ 4.5.   Proof of the technical lemma
$\mathop{}\!\mathbb{P}(\tau_{-r\log^{10K_{r}}(1/r)}>r^{2(1-1/\log\log(1/r))})$
$\qquad=\mathop{}\!\mathbb{P}(\inf\{W_{t},t\in[0,r^{2(1-1/\log\log(1/r))}]\}\geq-r\log^{10K_{r}}(1/r))$
$\qquad=\mathop{}\!\mathbb{P}(|W_{r^{2(1-1/\log\log(1/r))}}|\leq r\log^{10K_{r}}(1/r))$
$\qquad=\mathop{}\!\mathbb{P}(|W_{1}|\leq r^{1/\log\log(1/r)}\log^{10K_{r}}(1/r)).$
@@
For the last claim, to simplify notation set $r_{k}=r\log^{10k}(1/r)$. By symmetry, for any given $k$, the probability that $W$ reaches $-r_{k}$ before $r_{k}$ equals $1/2$ so if these events were independent, then this would follow e.g. from Hoeffding’s inequality. They are not independent, however by the previous bound, we may assume that at the first time $W$ reaches $-r_{k}$, it has not reached $r_{k+1}$ yet, and by the strong Markov property, the probability starting from this position to reach $-r_{k+1}$ before $r_{k+1}$ is (slightly) larger than $1/2$. Formally, one can e.g. apply the Azuma inequality to the martingale whose increments are
$\Delta M_{k}=\mathbf{1}_{\{\tau_{-r_{k+1}}<\tau_{r_{k+1}}\}}-\mathop{}\!\mathbb{P}(\tau_{-r_{k+1}}<\tau_{r_{k+1}}\mid(W_{t},t\leq\tau_{-r_{k}})),$
@@
(#$\mathop{}\!\mathbb{P}\Bigl{(}\#\{k\leq K_{r}\colon\tau_{-r_{k+1}}<\tau_{r_{k+1}}\}<\frac{K_{r}}{4}\Bigr{)}$ $=\mathop{}\!\mathbb{P}\Bigl{(}\sum_{k=1}^{K_{r}}\Bigl{(}\mathbf{1}_{\{\tau_{-r_{k+1}}<\tau_{r_{k+1}}\}}-\frac{1}{2}\Bigr{)}<-\frac{K_{r}}{4}\Bigr{)}$
$\leq\mathop{}\!\mathbb{P}\Bigl{(}\sum_{k=1}^{K_{r}}\Delta M_{k}<-\frac{K_{r}}{4}\Bigr{)}\leq\exp(-cK_{r}),$
@@
Similarly, for any $z>0$, the law of the time-reversed path $(z+W_{(\tau_{-z}-t)^{+}})_{t\geq 0}$ is that of a three-dimensional Bessel process started from 00, or equivalently the Euclidean norm of a three-dimensional Brownian motion, stopped at the last passage at $z$. The probability that a Bessel process does not reach $z/2$ before time $z^{2}$ is bounded below by some constant $c>0$ independent of $z$. We conclude as above, that the proportion of the number of indices $k\leq K_{r}$ such that this occurs for $z=r_{k}$ is bounded away from 00. ∎
@@
by a quantity that only depends on $r$ and whose sum over all values of $r\in\{2^{-n},n\geq 1\}$ is finite, such as $\log^{-2}(1/r)$. We then conclude as in the previous proof from the Borel–Cantelli lemma.
Step 1: Proof for Brownian motions. Let us first prove the claim when the Brownian bridges $W^{L}$ and $W^{R}$ are replaced by Brownian motions. Precisely, for every $k\leq K_{r}$, conditionally on $\Delta_{k}$, $\Delta_{k}^{L}$, $\Delta_{k}^{R}$, and $W_{G_{k}}$, let $X^{L,k}=(X^{L,k}_{t};t\in[0,\Delta_{k}^{L}])$ and $X^{R,k}=(X^{R,k}_{t};t\in[0,\Delta_{k}^{L}])$ be two independent Brownian motions started from $X^{L,k}_{0}=X^{R,k}_{0}=W_{G_{k}}>-r$.